OpenStudy (anonymous):

y varies directly as the square of R. if y=108 when R=6,find y when R=7??????

OpenStudy (anonymous):

y=kR^2

OpenStudy (anonymous):

uzma is right, 6^2=36 108/36=3 so k =3 when R is 7 y=3*7^2=147

OpenStudy (anonymous):

k=108/36=3 y=49*3

OpenStudy (anonymous):

y varies directly as the square of R. if y =16 when R=2 find y when R=7

OpenStudy (anonymous):

just put the values

OpenStudy (anonymous):

ok so what i have is y=kR^2 which would bring me out to R=7^2 which would be 49 which means y=k49 but then im lost I SUCK AT MATH AND i have a hard time comprehending how thigs work

OpenStudy (anonymous):

k=16/4=4 y=49*4

OpenStudy (anonymous):

ok i see now